Weekly Report for Week Ending December 8, 1998



 Project Control Meeting Agenda   Highlights   Hanford Observatory     Livingston Observatory     Detector    Project Office    System Engineering    Administrative    Past Weekly Reports

The Project Control Meeting Agenda for Monday  December 14, 1998 will be:
 (Meeting time: 10:30 am Pacific Time)

  1. Announcements (5 minutes)
  2. Review of Action Items (10 minutes)
  3. Comments on Weekly Report (5 minutes)
  4. Field Change Orders/Contingency Liens/Change Requests (10 minutes)
  5. Review of ISSUES AND PROBLEMS by Subsystem/Activity (10 minutes each)
Special Items:


Special Announcements:


Weekly Report Highlights
 
Footer problems with LHO staging building

First HAM airbearings at LHO, HAM8,9 get stacks, BSC pier alignment

Wrong fiber optic cable at LLO

Pictures from LLO
 
40 Meter recycling saga continues
 
Lightwave laser service experience
 

LLO bullet saga:

Note rods inserted in bullet holes point to location of shooter near LLO west end station

and note Dedham Springs, LA cartoon on the incident
 


LIGO Hanford Observatory (LHO) Operations



 General Items:
--------------
(Fred Raab)

Over the past week we hired several local high school students to help
out with many tasks around the observatory. So far, two are working out
well and one quit. C'est la vie! We also have a cooperative program
going with Mr. Hendricks' "Engineering Technology" class at Southridge
High in Kennewick. Student projects based on LIGO include a web page
overhaul, a virtual tour of LHO for the new web pages, and an
"interferometric attache".

Facilities:
-----------
O. Matherny

George Grant Construction Co. is building the forms for the concrete
footers and the large retaining wall for the staging building. Due to
lack of proper communications among the A/E consortium, the contractor
manufactured the footer forms and re-enforcing bars the wrong
dimension. This is causing some delay and possibly a monetary impact.

Apollo is 30% complete with water modification project. The tank
delivery is a long lead item. All other work is complete.

Bake Oven:
----------
(K. Ryan)

The RGA filament was destroyed earlier this week when the RGA
volume was vented while it was still hot. This was an "operator error"
on my part. LHO Vacuum Bake Oven A load #6 consisting of four 3" IOO
optics has completed its vacuum bake but remains unreleased as I have
not been able to perform the required post-bake scan. Replacement
filaments arrived and were installed late yesterday. A resistance check
of the newly installed filament indicates that it is not conducting.
Comparison with one of the other replacement filaments indicates that
the Thorium based coating sprayed on the filament wire during the
manufacturing process may need to be removed from the filament's
electrodes before the filament will work.

The removal of the RGA section for repair has provided an opportunity
to make design improvements in the choice and arrangement of the vacuum
fittings comprising the RGA volume. The previous arrangement had a
large mass of components hanging in free space at the end of what
amounted to a large lever arm connected to the next volume by a small
diameter flange. This arrangement was very intolerant to mechanical
loading (i.e. accidentally bumping into it) after being "thermally
exercised". Vacuum leaks have been induced in this way in the past. The
new arrangement cuts in half the length of the effective lever arm and
utilizes a much larger diameter connecting flange. Also, I have
fabricated a thermally insulative vertical support stand which, in
conjunction with the other changes, should eliminate this problem.
 

Seismic Isolation Installation:
-------------------------------
(C. Gray, H. Radkins)

At LHO:

Air Bearings were installed for HAM 7. This seemed to go smoothly.
Load within chamber was transfered from Support Brackets (Yellow
weldments bolted to Piers) to Actuation Stack (Air Bearing, Scissors
Table, Piers). This chamber needs only to have its remaining ~250 lbs
of weights installed and will then be complete.

Installation of In-vacuum components for HAMs 8&9 is complete (except
for all of 500lb payload in HAM 9 and ~250lbs in HAM 8). Doors will be
torqued down for both chambers today and diagonal section will be
pumped down thereafter. Installation went relatively smoothly; several
Leg Elements were rejected upon inspection for HAM8 because of excess
particles (most likely wood/sawdust ground up from within their
respective shipping crates) discovered.

A strong presence of visitors made substantial progress possible on the
BSC Pier alignment. Vertical elevation/level of Piers was completed on
several chambers. Horizontal alignment of chambers, utilizing "Big
Blue" fixture, was used at Y-outstation; initial results were not as
desired and will need to be looked at.

BSC 1ST ARTICLE

After installing mixed (right & left handed) springs into the isolation stacks on the 1st Article test
assembly in HYTEC's garage, instrumentation was calibrated and the performance testing began.
Using the PZT actuators driving in the U direction, response data was collected with two TAS
3000 3-axis geophones on both the support table and optics table in two configurations each.  The
SigLab measurement data collection system (Matlab based) allowed collection of all six channels
from the geophones and an input signal at the same time.  Using a random input signal and just four
setups, all degrees of freedom were collected with redundant data on all dofs except for pitch and
roll.  No V-motion response data was collected.
The PEM group's shaker/controllers were used to drive the table in the W motion with the same
data configuration collected.  The shakers were more limited in their input and useful frequency
range of data was more limited than the U driven data.  The results are still being worked through
but during collection no surprises were seen.

Optics & Lasers:
----------------
(D. Cook)

The LVEA particle counts continue to show large improvements. The
"tacky" mats and shoe covers are keeping a lot of dirt from being
tracked into the LVEA. The "tacky" mats last about three times longer
now that we are using shoe covers. We are now running the floor
scrubber twice a week throughout the LVEA. We have started to monitor
the dirty water that the floor scrubber vacuums back up into its
reservoir looking for signs of improvement. We are also having the
"hard-to-reach" areas under the beam tube, the beam tubes and chambers
wiped down, vacuumed and mopped once a month.

We will be switching the LVEA and VEAS into the "Laser Hazard"
condition approximately around the first week in January. This requires
all who enter these locations to be wearing laser safety eye wear. If
you plan on entering unescorted you will need to complete the site
laser safety requirements (LIGO-M980042-B-W). There are currently two
"Standard Operating Procedures" that are required reading, along with a
brief walk-through to answer questions and to point out things specific
to the laser installations. Please contact myself or Rick Savage for
further information.

Computing Systems:
------------------
(C. Patton)

Over the weekend the name of the server was changed from ligo168 to
rainier. This required saving the old nis+ database files to different
file names, re-initializing the nis+ software which created new
database files, then moving the old database files into the new
database files. I tried moving all the old files to new at one time and
found the server would not reboot. So I ended up having to move files
one at a time and reboot the server in between each move to find the
files preventing the server from rebooting. There are still some files
that need to be moved and I still have to figure out why some files
caused problems for the server.

All of the application software has been restarted with the new server
name after editing their license files. I won't have to re-license any
of them.

Electronics/Electrical Systems:
-------------------------------
(R. McCarthy)

Worked with the PSL team in characterizing Servos for the Laser. Work
went quite well. We met with Hytec to discuss cable routes for the
seismic racks. Josh took measurements for distances to BSC following
cable tray routes. Worked on misc. facility maintenance items, from bad
phones to well water pumping.

Beam-Tube Bake Out:
-------------------
(M. Lubinski, M. Guenther, and W. Althouse)

The bakeout of Y1 continues.  Tube temp is currently being controlled
at 160C. We have had some problems with PS1(DC power supply) and are
investigating the cause, but this has not slowed or stopped the bake
process. Some of the Cryopump cold plate temps are currently running
higher and we are investigating the cause.
 


LIGO Livingston Observatory (LLO) Operations


Vacuum Equipment: PSI has successfully completed the leak check of the x arm mid point gate valve and the man way used for entry into the BT for its repair. The x arm
manifold in the corner station is being baked. This is the last significant activity for PSI at LLO. They expect to be complete, except for punch list items, by the end of the year. (Sibley) Mike Zucker and Kyle Martini successfully completed seismic and acoustic measurements aimed at comparing the seismic noise environment with the specifications and understanding any anomalous effects.

Electrical Lab: The fiber optic cable being installed between the corner and end stations does not comply with the specified requirements in the contract (the calbe is only plastic jacketed and double armored cable is specified). We have stopped the installation work pending a response from the contractor describing how they intend to meet contract requirements. The racks in both end stations are being prepared so that the fiber optic cable can be installed. The racks need to be mounted to the ground and some patch cables for the network system connection need to be installed. Julien has been debugging CDS's vacuum control system with Russ Wooley from PSI
to start the bake-out of the LVEA-X arm manifold. We continue to buy more equipments to set up the EE lab. (Svoboda).

Optics & Vacuum Labs: Fisher/Hamilton has almost completed repainting of the tall cabinets, and will install them next Monday, December 14. The laser safety curtain has been received and as soon as Fisher/Hamilton installs the tall cabinets, Excel will hang the laser safety curtain, and install the laser interlock switch. Excel has also placed duplex outlets under the lab sinks, to power small water heaters for tap water. Culligan now expects to begin installation of the RO/DI water system late next week. We are still waiting for Woodrow Wilson to remove the four threshold plates from the doorways in the lab areas. To view the images, see: <ftp://ligo.caltech.edu/pub/Livingston/Misc/CableTrays.JPG

Facilities: Some nice new aerial photos of the Livingston site are now available. They were taken by a subcontractor to Hensel-Phelps, and they passed copies on to us. See: <ftp://ligo.caltech.edu/pub/Livingston/AerialPhotos/   Wye electric is continuing its work on the electrical installation cleanup. A HPC QA person is now assigned full time to oversee this work. We continue to chase cows off our property. Just how and where they get through the fence remains a mystery. A meeting on Thursday finalized negotiations between LIGO (A. Rizzi, Mark Coles) and LSU (Kevin Robbins, James Finney) for an industrial grade weather station to be operated and maintained by LSU weather scientists.The agreement is for one weather station at the corner station with the possibility open for additional units at each end station. (Rizzi).

Beam Tube Bake Out Preparation: DiCioccio Janitorial Service has been given a PO to pressure clean the Y-Arm Beam Tube & Enclosure. This work has been scheduled to start on 12-15-98. We have started installing "X -1arm" Beam Tube temperature sensors in their respective locations per LIGO document Ion LIGO-T970148xx. This work is schedule for completion by 12-25-98. Kerry Stiff measured the quality of the ground rods left behind by CBI and found them to have acceptable resistance to ground (about 8 ohms). A work order was issued to Excel to ground the beam tube at 28 places (at each double door). We continue to order Support Equip.and Consummable Goods. (Franklin, Stiff).

LSU related: A. Rizzi been appointed adjunct Professor at LSU. The LSU Security Office will visit the site Friday to view the bullet holes in our west end station and to suggest approaches for followup with the State Police and the FBI.

Administrative Issues: The passport software has installed and Bonnie has received the credit cards. The Christmas tree is up - thanks to everyone for their purchases of LIGO Livingston T-shirts! (Wascom)


WBS 1.2, 1.3    Detector Group


40 m Interferometer (Nergis Mavalvala)
---------------------------------------

Many tests have been done to understand the why the full recycled
interferometer falls out of lock when we align, and we are converging
-- though with some persisting ambiguities -- on the scenario that as
the ifo is aligned the optical gain in the RM and BS signals decreases
(to zero???) and lock is lost. The increased fluctuations in the better
aligned state are due to the decreasing gain until the gain is not
enough to keep the power recycled Michelson (PRM) locked.

To clean up our measurement of the RM optical gain (lock-in technique
described last week) we have implemented wavefront sensing for the RM
pitch/yaw degrees of freedom. This suppresses the large (>80%) 1 Hz
fluctuations by ~3x to the 30% level. With this reduction in 1 Hz
fluctuations, we confirm that the signal tends toward zero as we
align the ifo.

In an attempt to measure the locking behavior of the arm cavity, we
have introduced a second RF modulation frequency (fm2) which is near
but not at an integer times the fsr of an arm cavity. When the carrier
is exactly resonant in the arm cavity we would see no amplitude
modulation at fm2 in the transmitted light. However, if the cavity
is offset from perfect resonance for the carrier then a characteristic
transmission curve can be mapped as a function of fm2. The transmission
"dip" is wider for greater carrier detuning. We use this "SB probe" at
fixed fm2 to observe the carrier detuning in the arm as a function of
ifo alignment. The carrier does indeed detune further from resonance
(SB probe signal increases) as the ifo approaches better alignment.
The fluctuations in the SB probe signal are also reasonable: maxima
in probe signal coincide with minima in arm cavity power. This is
further evidence that as ifo is better aligned, it is being pulled
off resonance (one or more loops copping out).

Some conflicting evidence: (1) our measurements of visibility and arm
cavity losses indicate that carrier is well undercoupled; (2) we failed
to lock the ifo with good alignment using a dual sign flip on both the
RM and BS servos triggered on the zero crossing of the signal from the
lock-in (RM optical gain); (3) dual sign flip triggering on both arm
cavities' transmitted power also failed; (4) we can't compensate loss
of optical gain by increasing electronic gain in RM or BS servos beyond
a certain degree of alignment.
 
----------------------------------------------
Interferometer Sensing and Control -- M Zucker
(WBS 1.2.1.1.6, 1.2.1.1.7)
----------------------------------------------
 
Component fab/delivery status (MEZ)
-------------------------------------
item              status
----------------------------------------------------
IAS initial alignment tooling     delivered
WA2k&4k viewports      delivered
WA2k IO optical lever (MMT3)       delivered
WA2k COC optical lever optics    in fab, delivery 12/20
WA2k COC optlev supports (MMT3 type)    delivered
WA2k COC optlev supports (BSC type) in fab
WA2k COC optlev supports (TM type)  out for bid
WA2k video/illuminators        delivered
WA2k IOT7 WFS table        in final assembly, delivery 1/8 *
WA2k ISCT7,9,10          in layout, delivery 2-4/99 (scheduled)

    *Ready 12/15; shipping held to suit
    IO installation schedule.

ASC Design and Fabrication ( K.Mason, M. McInnis)
--------------------------------------------------------------------
Electrical feedthrus and blankoff plates for ISC table enclosures have
been designed, ordered, and recieved. In addition the special
shock-isolated shipping crate for IOT7 was recieved.

Completed the design for mounting the PLX retroreflector to the SEI
table during initial alignment. The PLX mount must now be detailed.

IAS design & fab (MEZ for M. Smith)
--------------------------------------------------------------------
Finished wedge adapters for all but two BSC-type optical lever supports
and sent out for fabrication.

Solicited quotes on designs for ITM/ETM -type optical lever support
structures.  Due to severe interface constraints (PSI piping, cable
trays, gate valve appendages, etc.) the structures are looking fairly
complex; we're doing some value engineering to try and reduce cost and
improve thermal/drift immunity.

ISC implementation (P. Fritschel)
------------------------------------------------------------------
o Finished determining the digital filter modules that need to
be implemented for the LSC controls, and gave these to Lori & Jay.

o Worked on setting up our two VME processors and one SUN on
a private subnetwork. Finally discovered that the second ethernet
card in the SUN was bad; a new one has just arrived and will be
installed.

o Rana & PF picked up the pace on finishing the assembly of
IOT7 (optical table containing the mode cleaner detection
electronics). The electro-optic shutter tested out ok. The plan
is to ship the table for arrival at Hanford in the first week
of January.

LSC Lock acquisition (Ware)
-------------------------------------------------------
Worked on modifying SMAC to give open loop system gain info.

ISC supervisory controls/GDS interface (E. Daw)
-------------------------------------------------------
Installed a second reflective memory card
in a single vme crate at mit in preparation
for timing tests on memory copies from one
card to the other.

Found out how to use the GDS timing functions
to read out the current time in nanoseconds.

Downloaded ROOT from CERN and started finding out to
call its graphics utilities from C code.

Got Sam Finns simulated data code in matlab
(thanks Sam) and started looking at it.
 

-------------------------------------------
Lasers and Optics -- J. Camp
(1.2.1.1.2, 1.2.1.1.3, 1.2.1.1.4,1.2.1.1.5)
--------------------------------------------

Prestabilized Laser -- Peter King
--------------------------------------------
Lightwave Electronics
=====================

- On 12/04, Glen Truong from Lightwave Electronics came down to Caltech to
perform a maintenance call on the 126 MOPA #103, the laser that was
originally shipped up for the 2k IFO PSL installation.  The laser suffered
from a noticeable drop in laser output power and beam quality.
Approximately 2700 hours of run-time was accumulated on the laser.

    The NPRO master oscillator was checked, since there was some
discrepancy between measurements (of things like laser crystal temperature
and pump diode current) obtained at Caltech and at Lightwave Electronics.
The reason for the discrepancy was found to be that Lightwave Electronics
obtained their readings from the laser power supply, whilst at Caltech we
obtained our measurements by measuring across certain pins on the 15-pin D
connector on the back of the laser.  With that resolved, the master
oscillator was set to the settings reported by Lightwave Electronics in
their test report.  It was agreed that in future both Lightwave Electronics
and Caltech would take measurements from the 15-pin D connector.

    What followed was a re-alignment of the optical train to the power
amplifier and an inspection of the components.  Two items in the optical
train were replaced; a thin film polarizer mount and a half-waveplate
rotation mount.  After the re-alignment, more pump power was measured going
into the power amplifier (about 470 mW) than when the laser was constructed
up at Lightwave Electronics (about 450 mW).  Despite this the output power
of the laser was lower than the 11 W reported previously.

    A window, the one in front of the double-pass mirror, of the power
amplifier was replaced, after a burn marks was observed on the inside of the
window.  This could not be noticed by a visible inspection because of the
design of the laser and was only noticed when the power amplifier was opened.

    Inspection of the power amplifier internal mirrors and the
double-pass mirror was then performed.  All the mirror surfaces were
cleaned and a new double-pass mirror was epoxied into position.  Despite
all these measures the output power of the laser was lower at just under 10 W.
The reasons for the power drop are not clear but must be related to the
condition of the power amplifier pump diodes.  The output power of the
laser was recovered only when the current to the power amplifier pump
diodes was increased by 1 A, which represents approximately a 5% increase.

    Performance of the laser will be monitored over the next few weeks.

Pre-stabilized Laser (PSL)
==========================

- Measurements characterizing the performance of the frequency stabilization
servo have been completed.  Measurements of the closed loop transfer
function indicate that the servo is performing as designed.  The servo
bandwidth was measured to be around 300 kHz.

    A new high voltage electro-optic modulator driver was tested.  The
unity gain frequency of the frequency stabilization servo with the new
driver was at least 900 kHz.  The increased unity gain frequency came at
the cost of dynamic range.  This was noticed by an increase in the
frequency noise whenever a transient event (such as closing the table
enclosure doors) occurred.  Performance was recovered by lowering the servo
gain and then increasing the gain to the previous level.  Whilst the higher
unity gain frequency was desirable, it came at the cost of robustness.  The
decision was made to stick with the old electro-optic modulator driver and
develop the new driver at a later stage.

- Measurements documenting the performance of the pre-modecleaner servo
have been completed.  Closed loop transfer functions and residual noise
spectra were obtained.  The closed loop transfer function indicated a servo
bandwidth of a few hundred hertz, which is consistent with the design.  The
residual frequency noise spectra indicate that the pre-modecleaner is
within specification for frequencies below 1 kHz and about a factor of 3
above specification for frequencies above 1 kHz.  This is consistent with
the performance obtained at Caltech and is a known limitation of the
current design pre-modecleaner.  This will be addressed sometime in the new
year.

- Both the frequency stabilization and pre-modecleaner servos have remained
locked overnight, a period of just over twelve hours.  Both servos are
still locked as of the time of writing this report.

- The IOO/PSL optical table surface has been grounded to the
instrumentation racks.

- Installation of the PSL intensity stabilization servo is about to
commence.  The acousto-optic modulator is installed and the position of the
output lens, which places a beam of the right size (0.5 mm) and location at
the IOO/PSL interface, has been marked.  Currently the workings of the
acousto-optic modulator driver are being studied.

- A new PSL EPICS control screen, which allows the operator to tell
at-a-glance the status of the PSL has been generated and tested.

- The status of the PSL construction for the other IFOs is unchanged from
the last reporting period except for the following:

    i. The installation of frequency servo and pre-modecleaner lock
acquisition software should be completed by 12/19 for the LHO 2k IFO.

    ii. Delivery of the new reference cavity vacuum chambers took place
last Thursday (12/03).  The new vacuum chambers will be scheduled for
cleaning and baking.

    iii, The Minco process controllers used in the thermal
stabilization servo are scheduled for shipping on 12/21.

    iv. The physical installation of the LHO 2k IFO PSL is expected to
be completed by 12/11.  The period 12/12 to 12/18 is earmarked for
performance and acceptance testing.
 
 

Input/Output Optics -- Dave Reitze, UF
-----------------------------------------

Task              Status
__________________________________________________
PSL IOO installation
    - optics on table   underway
    - periscope on table     complete
    - beam alignment     begins 12/19
    - beam characterization    begins 12/19
       - bench testing components underway

Vacuum Prep
    - suspensions       complete
    - nonsuspended optics  underway
    - suspended small optics   underway
       - riser block MMT2         to be shipped this week
    - suspended MMT3     begins early January

HAM7/8 installation
       - suspended optics (sos)   to begin 1/15
       - nonsuspended optics      to begin 1/15
 

IOO installation-LHO 2-k interferometer.

Qize Shu and Sany Yoshida have been at LHO all week. David Tanner will
replace Qize next Monday; Sany stays next week as well.

4 of the 3 inch mirrors are in the vacuum oven for bakeout after gluing
standoffs and magnets. The remaining 3 mirrors for the 2-k IFO are being
glued and balanced this week.

Bench testing of the electro-optic modulators is under way.

Parts fabrication--UF

The riser blocks for the MMT2 (to accomodate half the difference in heights
between the beam in the mode cleaner---set by the height of the SOS---and
the required height of MMT3) have been completed. The one for the 2-k
interferometer will be shipped to Caltech for baking this week.

The beam tube between the PSL/IOO enclosure and HAM 7 has been designed and
is in production.

Other

Work has been done to complete the list of the global positions of all
in-vacuum IOO components.

Work is beginning to measure the radius of curvature of MMT3 usng a cavity
scheme. A simple mount for the mirror has been fabricated, cavity paramters
calculated, and most of the apparatus (a laser, mirrors, mounts,
modulators,  detector, RF oscillator, and frequency counter) collected.
 
 

Core Optics Procurement and Metrology -- GariLynn Billingsley
-------------------------------------------------------------

We are taking final data on RM01, it will be shipped to Hanford December
28th.  This will be followed in Mid January by a 2ITM, we will check all
2ITM radii so that we can choose the best matching pair to work on first
for shipment.  All data are stored without reference subtracted so that we
can post process again any time we think we have a better understanding of
flat C.

Using the latest version of C (a combination of Fritz's method and the
CSIRO method) our difference  in units of waves from the CSIRO measurement
of side 2 of COC-A001 (HDOS) is:
    This week  Last week    Term
Z3  0.00215   0.002592    Power
Z4  0.00034   0.001070    Astigmatism at 0 degrees and power
Z5  0.00176   -0.002707   Astigmatism at 45 degrees and power
Z6  -0.00148 0.002598    Coma and X tilt
Z7  -0.00019 0.002545    Coma and Y tilt
Z8  0.00050   0.000837    Spherical Abberation and Power
Higher order term subtraction comes from the same source as last week, all
terms are ~< 0.5 milliwaves (yes, we're inventing our own units now!)
 
 

Core Optics Coatings -- Helena Armandula
--------------------------------------------------

Helena has been ill this week (JC)
 
 

Core Optics Support -- Michael Smith
------------------------------------

BEAM-DUMPS
The double beam dump prototype is pending assembly.

ETM and PO TELESCOPE
Completed sets of working drawings have been received for the ETM and PO
telescopes. Several details are still pending for the PO telescope. Ken M.
is evaluating the drawings.

VACUUM WINDOW
Jonathan is evaluating alternative approaches for the vacuum window, in
view of the excessively long lead time and the erratic quality reports on
the ISI windows.

PO MIRROR
The prototype PO mirror assembly will be evaluated next week when Jonathan
is at CIT.

NEW CONICAL BAFFLE at MID STATION
Design is pending.

TRIPLE FLEXURE BRACKET
Craig is working on a design of a triple flexure bracket (this was a
recommended change from the triple hinge concept at the FDR) for mounting
the beam dump horizontal support beams to the BSC chamber brackets.

MODE CLEANER BAFFLE
Mode cleaner baffle working drawing have been sent to a machine shop. The
glass is in the process of being cut by a vendor. The completed parts are
expected before 12/21/98. Cleaning and vacuum prep are expected to be
completed in time for installation in the LHO on 1/5/99.

IO BAFFLE
A top level design for the IO baffle, which will be mounted in HAM8, has
been completed by Angel. Paul K. will begin detail shop drawings.
Manufacturing is expected to begin on 12/15/98, with installation in LHO on
1/14/99.
 

INSTALLATION PLAN
A preliminary COS Manufacturing and Installation Plan has been completed
and is ready for review. Critical path manufacturing dates have been
identified. It is expected that we will meet the critical path installation
dates shown below.

item                                    order placed          install date
Mode cleaner baffle                       12/1/98     1/5/99
IO baffle                                 12/15/98(est)     1/14/99
 

--------------------------
Isolation Group -- M. Fine
--------------------------

Seismic Design (Mike Fine)
--------------------------------------
1. BSC Coarse Actuator Delivery:
Hytec is currently lining up all of the components that will be required
to deliver a new BSC actuation system to LHO on January 15th.  Hand
Precision is currently making several of the long lead items and is
aware of all of the hardware that will be required to meet this schedule.
The coarse flexural pivot assembly is the long lead item and is being
manufactured, but it will not be required in the assembly until the
system is ready to be driven.  This can be retrofitted into the BSC
after the completed assembly if required. The electronic components
for two complete racks have been ordered and are being received.
2. BSC In-vacuo Hardware:
Some minor changes to the BSC optics table have been approved that will
allow Allied to put some tooling hole in the surface of the table.  This
will eliminate only five optics mounting holes and is not envisioned to
cause any technical issues.

3. Coiled Spring Fabrication:
Pegasus has uncovered what they believe may be causing the difference in
spring performance observed between the left and right handed springs.
The left-handed springs have consistently been achieving better damping.
They believe this may be due to a difference in the spring follower tooling
that forms the spring against the mandrel.  They are making 12 springs with
modified tooling for us to test their theory.

4. Near-Term SEI Hardware Delivery Schedule:
16 HAM Bellows (Livingston)                                ~12/15/98
2 HAM support tubes and 1000 lb balance weights            12/21/98

Suspension -- (Janeen Hazel/Mark Barton/Mike Fine)
-------------------------------------------------------------------
1. Report From Janeen:
-Telecon with Brookfield's John Bergeron, Brookfield QA Dick Trunbull
and Bill Tyler and Irena Petrac on Tuesday, 12/8. Dick Turnbull had accepted
six dimensions on an inspection report for a D960133 structure that were
out of spec. Bill instructed them to always dilineate out of spec dimensions
as such.

-John Bergeron promised a schedule by the end of the day Tuesday. It was
faxed but LOS2 and LOS3 delivery data incomplete. He sent Excel version of
new schedule with this info today but we were unable to open it. He will
fax it tomorrow morning.

-I accepted shipment of the four D970571 height adapters along with the
two D970579 height adapters. Accepted for shipment one D960133 structure.

-First Recycling Mirror suspension structure due to be shipped 12/23.

-Bill Tyler and I will visit Brookfield Friday 12/18. Hopefully we can
visit the additional welding vendor at that time.

2. SUS Hardware Delivery:

SUS Structure Delivery Dates: [numbers in parentheses are quantities.]
LOS1(for ETM, ITM, MMT): 12/08(1), 12/18(2)*, 1/8(1)*, 1/22(1), 1/29(2),
                         2/5(2), 2/12(2), 2/19(2), 2/26(1)
LOS1(for RM): 12/23(2)*, 12/30(2)*
LOS2(for BS): 1st artcle on 1/15*, 2/15(2)*, 2/28(1)*
LOS3(for FM): 1/30(2)*, 2/28(1)*
SUS Height Adapter Delivery Dates:
MMT3, RM: 12/9(3)*, 12/11(3)*
FM: 12/9(1)*

SUS Fixtures and Components Delivery Dates:
Clamps: 12/14(20)*, 12/18*
BS Test Mass: 11/24*

* change

---------------------------------------------------------
Detector Systems Engineering -- D. Shoemaker/D Coyne
(WBS 1.2.1.1.9)
---------------------------------------------------------
 

Global Diagnostics System (Daniel Sigg)
--------------------------------------
1. John Z. prepared a design description for the intelligent
data monitoring/quick look diagnostics tool. It is accessible
at: http://www.ligo.caltech.edu/~jzweizig/DMT-Project.html

2. Had a discussion with Fred R., Rick S. and Haisheng R. about
data monitoring and analysis projects involving the laser and the
mode cleaner.

3. Work continued on the diagnostics test tool. Wrote the data
input module which will receive data from the network data server,
make a pre-analysis (decimation, heterodyne down-conversion,
decimation, partitioning) and saves the result into memory/disk
for final analysis.

4. Peter F. started writing the sine detection routine which will
do the following steps:
o as an option, remove the mean from the data
o mix it down at the detection frequency
o apply an FIR low-pass filter to the re & im parts of this
o apply the integration algorithm to these two data sets
o return a complex number describing the parameters of the sine.
 

Physics Environment Monitor (Alex Marin)
-------------------------------------------

Physics Environment Monitoring
--------------------------------
(Matt S., David. S, Alex M.)

    1.The new delivery schedule which is posted at :
       http://web.mit.edu/marin/www/PEM_Delivery.html
The changes are presented with italic bold characters.
    2.More work has been done for the detailed PEM installation plan, which
will be posted soon.
 

Optical Contamination (Daqun Li)
--------------------------------

13 cycles of vacuum venting and pumping for cavity one (with viton)
in two days do not result in a significant change in either ringdown
or mode spacing data.  Another set of viton has been requested so that
these tests may be repeated on a new batch.  Ringdown and
mode spacing data are still taken for cavities two (empty) and three (with
faraday).

Transmission map of a REO 3" flat optic (T=1500 ppm) has been obtained
using the automated Loss Scanner with a result of ~ 2% variation in
coating's transmissivity over the entire coating surface except the
edges where large transmissivity is observed as expected.  Transmission
mapping of a 10" optic may be needed for comparison.

 
-----------------------------------------
Control and Data System -- R. Bork
(WBS 1.2.2)
-------------------------------------------
No report received.
 


WBS 1.4    Project Office



 

 1.4.1.2 Project Controls (Lindquist)


Reports, Proposals, Work Plans (Lindquist)

The end of October Monthly Status Report is being finalized.

Materials for the end of FY 1998 Annual Work Plan are due by Friday, December 11, 1998.


Change Control (Lindquist)

The following Change Request has been submitted:
 
CR-980048 WBS 1.1.4 Extending Parsons Support in Livingston F. Asiri


LIGO Weekly Site Telecon (Jasnow)

The Site Telecon was held Thursday, December 10, 1998. The open action list is as follows:
 
Action No. Description Responsibility Assigned Date Due Date
1 Prepare Operations Budgets for FY 1999  Lindquist August 6, 1998 November 19, 1998
10 Procedure to Cover Late Charges for Invoice Payments Jasnow August 27, 1998 November 19, 1998
17 Louisiana Road Access M. Coles/G. Stapfer October 8, 1998 November 19, 1998
19 Consider Lawsuit Against Panel Manufacturer at Livingston R. Pool/E. Jasnow October 15, 1998 November 19, 1998
20 Schedule Meeting on Shadow Accounting System for Operations P. Lindquist October 15, 1998 November 12, 1998
21 Letterhead and Business Cards for Sites F. Raab/M. Coles November 5, 1998 November 12, 1998
22 Chart of Operations Off-campus Accounts for Staff P. Lindquist November 5, 1998 November 19, 1998

PROPERTY MANAGEMENT (Chargois)

>From: Ed Chargois <chargois_e@ligo.caltech.edu>

Assisted the Optical Engineer with the packaging and shipping of one (1) carton containing one inch Optics to be sent to the University Of Rochester, attention: A.C. Melissinos, scheduled arrival date of 12-9-98.

Assisted the Optical Engineer with the preparation of a Commercial Invoice and the shipment of four (4) containers with carriers SNs: 008, 009, 010, and 011 via Centra World Wide to Australia (CSIRO Division of Physics, attention: C.Walsh/A. Leistner.

Prepared and distributed SF form 122, Order NO. 99-NC-063 through the National Science Foundation (NSF) to the Defense Reutilization Marketing Office, Pensacola, Florida, to acquire a 10 inch Southbend Lathe, SN: 22849R with an acquisition cost of $12,000, to be shipped to LIGO Livingston Observatory. Transportation is being arranged through United Motor Freight.

Prepared and distributed SF form 122 Order NO. 99-NC-064 through the National Science Foundation (NSF) to the Defense Reutilization Marketing Office, Corpus Christi, Texas, to acquire an Engine Lathe (Monarch) SN: P7093 with an
acquisition cost of  $65,495.00 to be shipped to LIGO Livingston Observatory. Transportation is being arranged through United Motor Freight.


DOCUMENT CONTROL CENTER (Turner)

>From: the DCC <dcc@ligo.caltech.edu>

No report this week.


COST SCHEDULE CONTROL SYSTEMS (Duncan, Rianda, Leonhardt, Akutagawa, Kaufman)

>From: Cindy Akutagawa <cindy@acrux.ligo.caltech.edu>

I completed the worksheet and distribution for the MIT invoice for August and received the MIT invoice for September and October.  I have been doing various error checking and troubleshooting.  I am waiting for the CIT status reports for the month of November.

>From: "Britt Leonhardt (T)" <brittl@FINANCE.caltech.edu>

This week we have been busy coordinating with Florence, Irena, and others to iron out the Lightwave contract and account status.  I have spent some time going over the MIT invoices and the overhaul of the contract with Phyllis and Cindy.  I have handled more than the usual vendor calls regarding tax issues and faxing of tax-related documents to their
satisfaction.  I had a meeting with Katie from Price Waterhouse regarding the retention status of existing contracts and spent some time explaining their purpose.
 
WE Date Incoming Invoices and Receivers Wire Transfers Large Contract Invoices New Purchase Orders
10/08/98 248 2 9 155
10/15/98 225 0 4 157
10/22/98 247 0 1 58
10/29/98 182 1 4 42
11/5/98 141 2 20 41
11/12/98 132 1 11 63
11/19/98 118 2 8 14
11/25/98 61 1 4 4
12/3/98 186 2 15 102
12/10/98 175 0 6 90

>From: Kris Duncan <kris@ligo.caltech.edu>

Continued work on the November 1998 LIGO CSSR.  Analyzed and input all earned value and schedule status received to date.  Converted and posted the actual cost detail file from CIT finance.  Initiated a reconciliation of the LIGO vs. finance costs and commitments.

Continued to incorporate comments into the October NSF monthly report.

Documented the LIGO scheduling responsibilities to finish the construction effort.

>From: Florence Kaufman <fkaufman@ligo.caltech.edu>

Continued with cleanup of  IFIS records. Adjustments are being proposed for Beausoleil, Ray Capital Industries, D Hittle and Associates, Heritage Landscaping, Molding Solutions Incorporated.

Prepared a summary of the F-53 reports (commitments) by vendor and org (account) as of October 31, 1998 for distribution to the LIGO task managers.

Prepared an analysis of the amounts in the 3283 sub account (CIT-owned) as of the end of October.  The results show a total of $8 million, and ten subcontracts account for over 50 percent of the total. Further analysis is required to determine if these subcontracts have been correctly classified.

Assisted Accounts Payable in resolution of problem relating to the payment of the Lightwave invoice.

Provided auditors with information regarding retentions.


SUBCONTRACTS MANAGEMENT (Petrac, Jasnow)

>From: irena@ligo.caltech.edu (Irena Petrac)

>From: Ed Jasnow <jasnow@acrux.ligo.caltech.edu>

Ed Jasnow, Larry Jones, and Gerry Stapfer concluded all contract negotiations with CBI during a meeting held last week in Livingston.  The issues of extra leak checks and credits for LN2 dewars were settled, and the only item remaining to be added to the contract is $7,500 for shipment of some items from Plainfield to MIT.

It has been discovered that the fiber-optic cable company, MCA, has started to install un-armored cable in Livingston, in violation of the contract specifications.  He has been ordered to stop work, and is proposing the use of an aluminum sleeve to cover the cable.  Samples of this sleeve will be submitted for approval by LIGO before installation.

Just prior to pouring the footings for the new staging building at Hanford, the structural engineering subcontractor of the architect-engineer, D.H. Hittle, reported that the 5' x 5' footings approved by the architect, Dillman-Luvaas, in the shop drawings, would not be structurally sound for the building being erected.  The construction contractor, George Grant, has stopped work, awaiting revised drawings.  We are waiting for the delay claim from George Grant so we may know the exrent to which we will claim damages against the architect-engineer.


TECHNICAL CONFIGURATION MANAGEMENT (Althouse)

>From: wea@acrux (William Althouse)

No report. See Bake out.


1.4.2.1 Quality Assurance (Tyler)

>From: "William H. Tyler" <tyler_w@ligo.caltech.edu>

LIGO QA participated in the PSI Monthly Status telecon. There are two items that came up in conjunction with PSI  completing their work at Livingston that need LIGO attention. One item is to establish a spare/repair parts inventory for both sites which also addresses the fact that some items will not be available in the not too distant future. Another item requires a LIGO TRB to establish an acceptable in situ gate valve leak test of the replacement Flourel O-rings.

LIGO QA worked with Rita Torres to complete the draft of Mike Zydowicz's Site Operational Safety and Environmental Manual. Copies of this draft have been distributed for review and comment.
 


WBS 1.4.3    Systems Engineering / Data Analysis and Computing



 
 
Systems Engineering & Detector Integration Support

SEISMIC BALANCE WEIGHTS
    First shipment (1000#) has been received by the vendor who will be
cleaning & inspecting the balance weights. The second shipment (2000#)
should go out by Tuesday or Wednesday of next week. Subsequent shipments
will be going out when 1000# (pallet load) quantity is available to
allow cleaning vendor to proceed on a continuos basis rather than a lot
to lot basis. Material vendor ran out of stock so this job started
late but is now expected to come in ahead of schedule.

__Analyses: Nothing to report

__Reliability: Nothing to report

__Integrated Layout/Drawings:

INTEGRATED LAYOUT DRAWINGS (ILD) FOR LIVINGSTON
    Work has started on the ILDs for the Livingston site; however,
progress has stalled due to other priorities. Anticipate restarting
project
end of next week.

COVERS & GUARDS FOR HAM & BSC CHAMBERS
    Drawings are being updated for purchase of final quantities of large
& small O-ring guards & nozzle covers, for Hanford & Livingston. This
project is on hold for a couple of days to allow completion of baffle
drawings & ordering of hardware for early January installation.

X-BEAM & Y-BEAM SPOOL BAFFLES
   Detail drawings are in process but, low priority. New schedule is in
the process being generated

__Mock-ups: Nothing to report
 

Simulation & Modeling

* alfi

Some missing features and bugs of this releases were found and are
fixed. Send e-mail to emaros for bug reports. Please make the report as
detailed as possible.

* multi mode

The implementation of the summation cavity of the Michelson part is
going on.

* optics

The possibility of simulating the dual recycling configuration was
discussed. It is possible to do the simulation using the primitive
modules. In order to make it fast, we need to write a simple primitive
module, just like the summation cavity of Michelson.

* mechanical subsystems

S. Mohanty has reported that he is going to release a revised version of
his note, and will send us a matlab code of the single suspended mirror
soon.
G. Cella reported that the formaulations are all done, and expects to
have the first release integrated in e2e environment to be sometime in
February.

* LIGO-VIRGO collaboration in simulation/modeling area

We received an email from A. Vicere', who is in charge of the VIRGO
simulation now. He offered to start a collaboration between LIGO and
VIRGO. We had dicusssions internally, and the following are the
items to be considered.

(1) we need to have a meeting as soon as possible - possibly at STAIC II
    (a) understand the situations and requirements of each group's
efforts
    (b) understand the software program structure
(2) topics for consideration for collaborative work
    (a) validations
    (b) modeling of (how to implement) physics
    (c) computation - simulation speed up
    (d) sharing code development

Andrea is talking to the VIRGO simulation group about this
collaboration,
and I welcome any comments or suggestions regarding this effort.

* IOO

S.Klimenko is going to come to CIT at the beginning of January.
Biplab has explained the details of e2e to S.Klimenko over phone.

Data Analysis

A. Lazzarini: We held a joint meeting with the LSC chairs of the data
groups last Friday. This will become a monthly ritual from now on. The
minutes for the meeting are available at

http://docuserv.ligo.caltech.edu/~prince/LDCG_lsc/Emails/index.html

Worked with W. Althouse on revision of the BT database access program
via the www. This effort is almost complete as far as the temperature DB
is concerned. We stil need to address integration of the RGA partial
pressure DB.

B. Sears:

* Electronic Log software - making a more portable, maintainable,
        and customizable electronic log out of the FLNL's original
version.

Kent Blackburn:

* Met with Franham to discuss several minor modifications
  to the internal light-weight data format to support
  collections of XML elements.
 
* Met with Ehrens to discuss syntax for collections of LDAS
  commands being handled by the ManagerAPI queues.
 
* Worked with Hu to try and trace a DB2 client-to-server
  communication error. The error appears to be do to an
  IP identity typo on a nameserver somewhere. The problem
  was intermittent at first and seems to have resolved itself
  at this time.

Dave Farnham:

* I am continuing to work on the FrameAPI.  I project that
  the coding will be complete by 12/18.
 
* Spent some understanding the Fcl code and how best to
  implement the needed coding standards for the LDAS.
 
* Continued to probe ALTA about getting a new beowulf node
  to replace the defective unit.

Phil Ehrens:

* Completed about 30% of the managerAPI.tcl script.

* Developed coding standards for Tcl/Tk for which
  the TclDOC Perl script can be used as a validator.

Xiao Hu:

* DB2 installation and testing

* Loading data to a sample table according to Roy's suggestion.
  (before the SDK come, this can only be done by script)
 
* Reading document about generic API and manager API, get myself
  familiar with Phil and David's work. (Kent's suggestion)
 
* Studying the methods to accessing DB2 using Tcl/TK and the method
  to enable web-based querying using IBM product Net.Data.
 
* Making arrangement to get DB2 Appilication Developer's Kit.(returning
  Workgroup Edition in exchange for a Developer's Edtion).

General Computing (Wallace is out of town this week)

- Hanford --

(C. Patton)
1.  Over the weekend the name of the server was changed to rainier. The
nis  had to be re-initialized and the old database files moved over.
There
are still some database files that will be moved this weekend. All the
application software licenses were able to be restarted, after editing
license files, with the new name.

2.  All LHO computers are now on the DOE ESnet T1 line. Our domain name
is ligo-wa.caltech.edu. Our server is named rainier.ligo-wa.caltech.edu.

3.  Helped with multiple software problems on PCs. Created several new
user accounts. Followed up on several software orders. Worked on
reconciling and updating '98 and '99 expenditures spreadsheets.

4. [AL Note] Fred was able to gain access to Caltech's library services
at Hanford with the assistance of Caroline Smith from Library Services.
They now have (or soon will have) TOC/DOC, SCIsearch, etc. This became
possible once the LIGO WAN was set up. Similar services will be
available to the LLO staff.

-- Caltech --

(S. Singh)
1. The three printers (hpcolor3, hp4mlrtsn, lw3) are put online after
they  underwent necessary maintenance.

2. A couple of new user accounts have been set up. Some PC's problems
were
fixed.

3. Chip Sumner's (grad student) PC has been upgraded to 96 MB of memory
to
cope
up with memory intensive applications.

4. Purchase order has been place with bookstore for three Sun mice.

5. Fax quotation for 8mm 7-14 GB tape drive has been obtained from Sun.
Similarly, quotation for Adobe Framemaker 5.5 for Unix has been obtained
from
bookstore.

(B. Kratochwill)
1.  Installed the Amaldi Conference web site.  Backed up the database
and
web pages.  Wrote a paper report of applicants.

2.  Drew up a sample page for the as-built drawings.  Page is being
reviewed by F. Asiri.

3.  Installed photo page from October NSF meeting on internal bulletin
board.  Made several quick web site changes -- added Simulation bulletin
board,  several 98 publications, another LIGO talk.

-- MIT --

No report

-- Livingston --

No report
 


WBS 1.4.4.1    Administrative Group


 Rita Torres

For I. Petrac did change order No. 5 to University of Florida, change order No. 3 to Allied Engineering, change order No. 10 to National Institute of Standards. Distributed change order No. 8 to Butler Services Group.

For W. Tyler, did edits to draft of LIGO Safety and Environmental Protection Manual (LIGO-M980242-00-M). Distributed manual for review and comment, the document is now 57 pages. Also went over document list to compare with the DCC version; to verify whether documents had correct revision letters and dates. Also identified documents still needing work.

Collected orders from 2nd and 3rd floor folks for desk calendars. Ordered calendars and office supplies in one swoop.
 

Irene Baldon

Worked on the ever present backlog of paper work for various trips taken recently (40 Purchase Requisitions and 17 Advance Requests). Worked on the ever present backlog of Expense Reports (53 Expense Reports). The backlog is slowly but surely going down.

I am now arranging and preparing the paperwork for MIT trips to our 2 sites and CIT. Your cooperation in assisting me with the taping of your receipts on 8-1/2 x 11 white paper and filling out your own Expense Reports, will greatly help in processing your reports in a timely manner. This will enable me to do a speedier audit and only require changes when necessary. If you are not sure of the per diem, or anything else, please leave it blank, as well as the totals at the bottom, and I will fill it in.

Performed miscellaneous duties as requested by various members of the LIGO Project here at Caltech, as well as for members of the staffs of each of the two sites.
 

Dorothy Lloyd

Processed heavy load of purchase requisitions, POs, change orders, check requests, etc. as well as heavy load of invoices (see Cost Schedule Control Systems report submitted by Britt Leonhardt for further details).

Followed up and tracked problems with invoice payments.

Continued reconciling June status reports.

Continued to monitor and support the temporary help.
 

Elizabeth K. Wood

Just a reminder, I am doing payroll now. When you take a vacation or are sick, I’m the one to whom you need to send the pink sheets.

Sold many sweatshirts and t-shirts for the Livingston LIGO Employees’ Recreation Club.

Continued interviews with Syd Meshkov in the search for an assistant to help with the Amaldi Conference.

Had an extensive meeting with Alan Rice to update him on life in LIGO and the changes Caltech is facing in the coming months.

Met with Irene and Gary concerning many travel-related issues.

Updated the LIGO space assignments. Please keep me notified if you are expecting visitors so I can arrange a place for them.


For additional information about this report, contact sanders@ligo.caltech.edu